On September 30, 2021, the San Fillipo Corporation issued 8% stated rate bonds with a face amount of $280 million. The bonds mature on September 30, 2041 (20 years). The market rate of interest for similar bonds was 10%. Interest is paid semiannually on March 31 and September 30. ((EV of $1. PV of $1. FVA of $1. PVA of $1. EVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.) 

Required: 

Determine the price of the bonds on September 30, 2021. (Enter your answers in whole dollars. Round your final answers to nearest whole dollar amount.) 
Answer #1

Answer:

Calculate present value of face value of bond
Particulars Amount($)
Face value of bonds                  28,00,00,000
Pv factor at 5% for 40 periods 0.14205
Present value of face value of bonds                    3,97,74,000
Present value of $1 annuity table at 40 periods at 5% is 0.14205
Calculate present value of interest payment ( ordinary annuity table)
Interest payment (2800000000*8%/2)                    1,12,00,000
Pv factor at 5% for 40 periods 17.15909
Present value of interest payment                  19,21,81,808
Price of bond                  23,19,55,808
Answer #2
table values are based on
n= 40
i= 5%
cash flow Amount PV
interest 11,200,000 192181808
principal 280,000,000 39774000
price of bonds 231955808
(note there will be slight difference in issue price calculation ,kindly use the factors as given
inn your question to get exact answer)
Bond characterstics Amount
1-a) Principal 280,000,000
interest      11,200,000
Market interest rate 5%
periods to maturity 40
issue price 231,955,808
Calculation of bond issue price
Where
i= 5.00%
t= 40
principal * PV of $1 at 5% for 40 yrs =
280,000,000 * 0.14205        = 39774000
interest * PV of ordinary annuity at 5%=
11200000 * 17.15909 = 192181808
bond issue price 231955808